OpenStudy (anonymous):

True or False: Only polynomials with odd-numbered degrees have remainders. Select one: True False

OpenStudy (solomonzelman):

Something that can be false sometimes is false.

OpenStudy (anonymous):

False

OpenStudy (puzzler7):

You could have something like \[\frac{x^{2}+8}{x^{2}+7}\]

OpenStudy (anonymous):

I don't think that's a polynomial^

OpenStudy (solomonzelman):

When x→∞ I would approximate it to be 1 ;)

OpenStudy (solomonzelman):

yes, it is not a plynomial

OpenStudy (nincompoop):

zero

OpenStudy (solomonzelman):

??

OpenStudy (puzzler7):

@SolomonZelman that would be\[\frac{\infty}{\infty}\]which would be undefined.

OpenStudy (nincompoop):

it actually isn't undefined

OpenStudy (solomonzelman):

No when x approaches infinity, for any real number x, it would be ≈1 Also nincompoop points out, that ∞/∞ is not undefined (on some level)

OpenStudy (solomonzelman):

When you say x→∞ you don't actually mean that it is ∞ / ∞

OpenStudy (puzzler7):

You're right. Sorry, I solved it as x = infinity. Using L'Hopital's rule, it equals 1.

OpenStudy (solomonzelman):

Yes L'H'S would give 1, but what does it have to do with anything here?

OpenStudy (solomonzelman):

Remainder and L'H'S aren't so much related to each other... or at least not in a way that I aware of.

OpenStudy (puzzler7):

No, this had nothing to do with the problem. I was merely noting my stupidity and acknowledging that you are correct.